How many children have heights between 98 cm and 104 cm?

1
Expert's answer
2022-02-28T11:17:08-0500

Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. How many children belong to the upper 15% of the group?

Let "X=" height of a child: "X\sim N(\mu, \sigma^2)."

Given "\mu=110\ cm, \sigma=6\ cm."


"P(98<X<104)=P(X<104)-P(X\le98)"

"=P(Z<\dfrac{104-110}{6})-P(Z\le\dfrac{98-110}{6})"

"=P(Z<-1)-P(Z\le -2)"

"\approx0.158655-0.022750=0.135905"

"0.135905\cdot 500=68"

68 children have heights between 98 cm and 104 cm.